# Perception-as-Control: Fine-grained Controllable Image Animation with 3D-aware Motion Representation

### v1

@@ Abstract
Motion-controllable image animation is a fundamental task with a wide range of potential applications. Recent works have made progress in controlling camera or object motion via various motion representations, while they still struggle to support collaborative camera and object motion control with adaptive control granularity. To this end, we introduce 3D-aware motion representation and propose an image animation framework, called Perception-as-Control, to achieve fine-grained collaborative motion control. Specifically, we construct 3D-aware motion representation from a reference image, manipulate it based on interpreted user intentions, and perceive it from different viewpoints. In this way, camera and object motions are transformed into intuitive, consistent visual changes. Then, the proposed framework leverages the perception results as motion control signals, enabling it to support various motion-related video synthesis tasks in a unified and flexible way. Experiments demonstrate the superiority of the proposed framework. For more details and qualitative results, please refer to our project webpage: Perception-as-Control.
Figure 1: Potential applications for Perception-as-Control. By constructing 3D-aware motion representation based on user intentions and utilizing the perception results as motion control signals, the proposed fine-grained motion-controllable image animation framework can be applied to various motion-related video synthesis tasks, such as image-based motion generation (animate image according to user inputs), and video-based motion clone (mimic the entire motions), motion transfer (relocate and scale local motions based on semantic correspondence), motion editing (editing any motion in user-specified regions).
@@ 1 Introduction
With the proliferation of generative AI, image animation as a fundamental task is undergoing rapid progress [20, 7, 2, 40, 23, 16, 27, 31]. Various potential application scenarios put forward higher demands for this task, i.e., particularly in achieving precise motion-controllable image animation based on the interpretation of user intentions. Motion-controllable image animation usually takes a single reference image and motion control signals derived from motion representation as input, aiming to generate videos with the expected motions while maintaining the specific appearance. Since rich appearance information is already included in the reference image, the key lies in designing better motion representations and effectively guiding the motion content generation process.
@@
Motion, in general, can be decomposed into camera motion and object motion. Object motion refers to dynamic movements relative to the world coordinate system, while camera motion arises from viewpoint changes. Although researchers have made great progress in designing various motion representations to obtain control signals for camera or object motion control, achieving collaborative camera and object motion control with fine control granularity remains non-trivial. For camera motion control, some methods directly use camera intrinsic and extrinsic parameters [33] or convert them into Plücker coordinates [37, 9, 47] as high-level control signals, while others apply point cloud reconstruction to render incomplete frames [3, 43, 10] or use dense optical flow [23, 15] as more concrete control signals. For object motion control, recent works usually follow the drag-based interaction setting that allows users to draw point trajectories on the reference frame. These methods convert sparse point trajectories into optical flow [33, 16, 45, 15, 23], reorganize point embeddings [27, 34, 5, 4] as control signals. Nevertheless, works [15, 4, 33, 16, 5] focusing on collaborative camera and object motion control still suffer from conflict control issues and insufficient control granularity.
@@
The key observations of this paper are three-fold: 1) Motion representations based on pixel offsets such as optical flow are flexible to modify, while the lack of 3D awareness leads to conflict control issues in local regions with simultaneous camera and object motions. 2) Reconstruction-based motion representations such as point clouds show better spatial consistency, making them effective for camera motion control but less flexible for fine-grained object motion control. 3) For controllable visual content generation tasks that maintain basic spatial structures, diffusion models prefer spatially aligned control signals over high-level ones. Based on these observations, an effective motion representation for fine-grained collaborative motion control should have the following properties. First, the motion representation is equipped with 3D awareness to disentangle local components from the reference image. Second, the motion representation is consistent in space for stable camera motion control and flexible enough for fine-grained object motion control. Third, the motion control signals derived from motion representation are spatially aligned with the generated frames to enhance the connections between them.
@@
Figure 2: 3D-aware motion representation for simplifying and summarizing the 3D scene corresponding to the reference image, which consists of several unit spheres denoting key object parts and a world envelope accommodating them. We move those spheres and observe the scene based on user intentions to transform camera and object motions into intuitive and consistent visual changes.
@@
To this end, we introduce 3D-aware motion representation to facilitate fine-grained collaborative motion control. In controllable image animation applications, users generally have clear intentions about how objects move and how to observe the scene. The intuitive approach is to reconstruct the 3D scene based on the given reference image, drive the objects according to various user intentions, and render the scene from the expected viewpoints. Considering the significant effort required to reconstruct everything in the scene, we notice that human visual attention is predominantly drawn to moving parts, while the remaining parts are generally perceived as a whole [6]. Therefore, as shown in Fig. 2, we simplify and summarize the 3D scene as a combination of several unit spheres representing key object parts (most moving parts and a few static parts) and a vast world envelope akin to The Truman Show to demonstrate how the scene evolves as the observer’s viewpoint changes, i.e., 3D-aware motion representation. We then render the instantiated 3D scene to obtain perception results that reflect camera and object motions and necessary spatial relations.
@@
Based on 3D-aware motion representation, we propose a fine-grained motion-controllable image animation framework, called Perception-as-Control, to achieve collaborative control of arbitrary and fine-grained camera and object motions. The framework takes a single reference image and the perception results of the corresponding 3D-aware motion representation as motion control signals, which are spatially aligned with the reference image. Specifically, we apply two lightweight encoders with similar architectures to encode camera and object control signals separately to avoid RGB-level interference. The encoded control signals are then merged to form the final motion control signals. We employ a diffusion model based on U-Net architecture and inject appearance information from the reference image and motion information from motion control signals, enabling controllable image animation. We introduce a data curation pipeline to prepare training data from in-the-wild videos. Moreover, we propose a three-stage training strategy to balance camera and object motion control and facilitate fine-grained motion control.
@@
  • We introduce 3D-aware motion representation to facilitate fine-grained collaborative motion control, which can be modified and perceived to transform user intentions as spatially-aligned motion control signals.
  • Based on 3D-aware motion representation, we propose a fine-grained motion-controllable image animation framework, which takes advantage of diffusion models for achieving precise and flexible collaborative motion control.
  • The proposed framework can support various motion-related applications via user intention interpretation, such as motion generation, motion clone, motion transfer, and motion editing.
@@ 2 Related Work
Camera Motion Control.
@@
Works targeting on camera motion control [33, 37, 47, 9, 3, 43, 7, 29] aim to generate novel view images with spatial consistency. In computer graphics, camera information is typically represented using a set of intrinsic and extrinsic parameters. Methods like MotionCtrl [33] directly repeat these flattened camera parameters to image size as control signals. CameraCtrl [9] and CamCo [37] transform the camera parameters into Plücker coordinates as pixel-wise embedding, resulting in more discriminative camera motion representations. ViewCrafter [43] and I2V-Camera [3] lift the reference image to 3D and render the point cloud to generate incomplete frames. SparseCtrl [7] and DimensionX [29] generate a specific set of training data for each type of camera motion and train individual LoRA models [11] tailored for each camera motion pattern. Concurrent work TrajectoryAttention [36] uses dense pixel trajectories to enhance temporal consistency, which is effective for camera motion control but hard to further support object motion control. Compared to high-level camera motion representations, our 3D-aware motion representation embodies camera motion information via visual changes of the world envelope, which is more intuitive and effective.
Object Motion Control.
@@
Previous works usually follow the drag-based interaction setting for object motion control [42, 16, 27, 23, 45, 34], which conveys user intentions via sparse point trajectories. Methods such as DragNUWA [42], ImageConductor [16], Motion-I2V [27] and MOFA-Video [23] employ optical flow as an intermediate motion representation. The former two methods transform point trajectories into sparse optical flow as control signals, while the latter two methods transform sparse point trajectories into dense optical flows via off-the-shelf or pretrained sparse-to-dense flow generation models. Tora [45] enhances point trajectories into a trajectory map and employs 3D VAE to encode it. DragAnyhting [34] obtains features for each pixel via DDIM inversion [28] and reorganizes those features according to segmentation masks and point trajectories while being less effective in handling motions containing shape variations. Although the above methods support camera control using point trajectories, they suffer from conflict control issues in local regions with simultaneous camera and object motions. Instead of modifying 2D motion representation, we employ a simplified 3D scene and use its inherent perception results as control signals to avoid control conflicts. Due to this design, diffusion models can handle both local object motion and global camera motion simultaneously.
Collaborative Motion Control.
@@
Only a few concurrent works [15, 5, 4] support collaborative motion control. AnimateAnything [15] applies two-stage generation, one for transforming camera and object control into optical flow representation, the other for generating videos based on that. However, they represent camera motion as Plücker coordinates, which leads to imprecise camera motion control. MotionPrompting [5] employs dense point embeddings constructed from 2D point tracking results as motion control signals, while such representation is not friendly enough for complicated motion control. I2VControl [4] also employs dense point embeddings constructed via point transform and projection, while the granularity of object motion control is limited to object segmentation masks. In contrast, we propose Perception-as-Control, which takes motion control signals derived from efficiently constructed 3D-aware motion representation and achieves fine-grained collaborative motion control.
@@ 3 Perception-as-Control
Figure 3: Overview. In our data curation pipeline, an in-the-wild video is taken as input, and off-the-shelf visual odometry and 3D point tracking algorithms are applied to obtain camera poses and 3D point locations. Based on these results, 3D-aware motion representation can be constructed and rendered to derive camera and object control signals. During training, those control signals are fed into separate encoders to avoid RGB-level interference and merged afterward. A three-stage training strategy is introduced to facilitate collaborative motion control. During inference, user intentions are first interpreted as 3D-aware motion representation, enabling our framework to support various motion-related applications.
In this section, we first introduce a data curation pipeline for constructing 3D-aware motion representation from in-the-wild videos. Then we describe our fine-grained motion-controllable image animation framework, Perception-as-Control. A three-stage training strategy is introduced to balance camera and object motion control and facilitate fine-grained object motion control.
@@ 3.1 Preliminaries
Video Diffusion Model (VDM) adopts the foundational principles of image diffusion probabilistic models, leveraging architectures like denoising U-Net [2, 8] and DiT [40] for video generation. In addition to generating reasonable images, VDM aims to model the temporal dependencies between frames, generating videos that maintain consistent visual content and coherent motion. Let $x_{0}\in\mathbb{R}^{T\times H\times W\times C}$ denote the initial video latent, where $T$ is the number of frames to generate. The forward process of diffusion models is defined as a chain that progressively adds Gaussian noise to the original video, as shown in Eq. 1:
$x_{t}=\sqrt{\hat{\alpha_{t}}x_{t-1}}+\sqrt{{1-\hat{\alpha_{t}}}\epsilon},\epsilon\sim N(0,1),$ (1)
where $t\in{1,\dots,T}$ denotes the diffusion timestep, $\hat{\alpha_{t}}$ regulates the intensity of noise added at each $t$, and $\epsilon$ is drawn from standard Gaussian noise. Thus, the optimization objective of a diffusion model is to predict the probability distribution $p(z_{t-1}|z_{t})$ using a parameterized network $\theta$. The corresponding loss function can be expressed as Eq. 2:
@@
where $\mathcal{C}$ is an optional parameter denoting the conditions, such as reference images and motion control signals. Typically, VDMs first encode videos into a latent space with the same dimensions as $x_{t}$ to reduce the computational complexity and achieve better results.
@@
For training purposes, pairs of video data and the corresponding 3D-aware motion representation are required. Unfortunately, commonly used large-scale video datasets predominantly consist of in-the-wild videos lacking camera settings and scene information. One way to recover such information involves estimating camera parameters and reconstructing the scene through monocular 3D reconstruction algorithms such as SfM [46] and then simplifying the reconstructed 3D scene. However, those algorithms are computationally expensive and the scale ambiguity issue leads to non-negligible reconstruction errors. Alternatively, constructing the simplified 3D scene directly via appropriate approximations can bypass these problems and is readily achievable even for in-the-wild videos. Thus, we introduce our data curation pipeline, using off-the-shelf algorithms to construct 3D-aware motion representation for any collectible video, as shown in the top-left part of Fig. 3.
Cut Detection and Motion Filtering.
@@
First, we apply common preprocessing steps to extract video clips containing continuous visual content. We employ a cut detection algorithm and split the original videos into several video clips. Subsequently, we use RAFT [30] to estimate the optical flow between each pair of adjacent frames in a video clip. We then calculate the Frobenius norm of the estimated optical flow as a motion score, which is further used to filter out video clips with non-obvious motion (motion scores lower than $30^{\rm th}$ percentile), the same as previous methods [16]. Next, we construct our 3D-aware motion representation for each video clip. The process is divided into two steps: 3D point tracking to capture local object motion and visual odometry to capture global camera motion.
3D Point Tracking for Local Object Motion.
@@
Key object parts refer to not only moving object parts but also stationary ones, which are critical to revealing spatial relations. They are simplified as unit spheres in the 3D scene, and the number of these spheres is defined according to control granularity flexibly. We require 3D locations of key object parts at each moment to set these spheres in the initial location and simulate their movements. Therefore, we employ SpaTracker [35] to track 2D points in the pixel plane while estimating their depth relative to the camera, i.e., 3D point tracking in each video clip. Note that combining a 2D point tracking algorithm [13] with a depth estimation algorithm [24, 38, 39] has similar functions. In our implementation, we track $25\times 25$ grid points in each video clip, resulting in 3D point tracking results $P=[P^{\rm X},P^{\rm Y},P^{\rm D}]\in\mathbb{R}^{T\times N\times 3}$ ($T$ is the number of frames and $N$ is the number of tracked points).
Visual Odometry for Global Camera Motion.
@@
We employ a world envelope to mark the remaining parts, which is stationary relative to the world coordinate system. To transform global camera motion into visual changes, we can use off-the-shelf algorithms [32, 46] to recover camera trajectories from a sequence of continuous frames and render the world envelope. Despite the scale ambiguity problem, our world envelope, which serves as a scene marker, only requires coarse camera poses instead of extremely precise ones. Thus, we set the camera intrinsic parameters as $I=[W,0,W//2;0,H,H//2;0,0,1]$, where $W$ and $H$ are the width and height of the first frame. In our implementation, we employ TartanVO [32] for this purpose, since it is fast and generalizes well across various cameras. For each video clip, we obtain the estimated camera poses $E=\{E_{t}|t\in\{1,2\dots,T\}\}$, where $T$ is the number of frames, $E_{t}=[R\in\mathbb{R}^{3\times 3},T\in\mathbb{R}^{1\times 3}]$ are camera extrinsic parameters.
Perceptual 3D Scene Rendering.
@@
Once we have obtained the 3D point tracking results $P$ and the estimated camera poses $E$, we construct our 3D-aware motion representation and render it as motion control signals.
@@
First, we set $N$ unit spheres with different colors representing key object parts. These spheres are rendered on the pixel plane according to the camera pose at each timestep and appear as circles that are larger near and smaller far away in the rendering images under perspective projection. Since the estimated depth is not precise enough to determine the absolute 3D positions of these spheres, we normalize point depths $P^{\rm D}$ as $\hat{P}_{\rm d}$ for each video clip to only focus on their relative distance. After that, we set the projected circle radius of the $n^{\rm th}$ sphere to $1-P^{\rm D}_{t,n}$ ($P^{\rm D}_{t,n}$ is the normalized depth of the $n^{\rm th}$ sphere at timestep $t$). For convenience, we generate a continuous color map where each pixel has a distinct color. We then set the color of each sphere to the corresponding color of its projection center in the first frame. Finally, these spheres are projected and rendered as colorful circles with varying radii in the pixel plane.
@@
Then, we construct a cube centered at the origin with the side length of $z_{far}$ ($z_{far}$ is the far clipping plane of a viewing frustum) as an implementation of the world envelope. We apply UV textures in checkerboard style to increase the recognizability of the world envelope, which helps demonstrate how the scene evolves as the viewpoint changes. We align the camera coordinate system to the world coordinate system for the first frame and render the subsequent frames according to the camera pose sequence $E$, without loss of generality. Note that the implementation of the world envelope is not limited to a cube, and its textures can vary beyond the checkerboard style.
To avoid RGB-level interference, we render unit spheres and world envelope separately as two spatial layers, acting as object and camera control signals respectively.

### v2

@@ Abstract
Motion-controllable image animation is a fundamental task with a wide range of potential applications. Recent works have made progress in controlling camera or object motion via various motion representations, while they still struggle to support collaborative camera and object motion control with adaptive control granularity. To this end, we introduce 3D-aware motion representation and propose an image animation framework, called Perception-as-Control, to achieve fine-grained collaborative motion control. Specifically, we construct 3D-aware motion representation from a reference image, manipulate it based on interpreted user instructions, and perceive it from different viewpoints. In this way, camera and object motions are transformed into intuitive and consistent visual changes. Then, our framework leverages the perception results as motion control signals, enabling it to support various motion-related video synthesis tasks in a unified and flexible way. Experiments demonstrate the superiority of the proposed approach. For more details and qualitative results, please refer to our anonymous project webpage: Perception-as-Control.
Figure 1: Potential applications for Perception-as-Control. By constructing 3D-aware motion representation based on user intentions and utilizing the perception results as motion control signals, the proposed fine-grained motion-controllable image animation framework can be applied to various motion-related video synthesis tasks, such as image-based motion generation (animate image according to user instructions), and video-based motion clone (mimic the entire motions), motion transfer (relocate and rescale local motions based on semantic correspondence), local motion editing (edit fine-grained scene and object motions in user-specified regions).
@@ 1 Introduction
With the proliferation of generative AI, image animation as a fundamental task is undergoing rapid progress [22, 8, 2, 44, 25, 18, 29, 35]. Various potential application scenarios put forward higher demands for this task, i.e., particularly in achieving precise motion-controllable image animation based on user intentions. Motion-controllable image animation usually takes a reference image and motion control signals derived from motion representation as input, aiming to generate videos with the expected motions while maintaining the specific appearance. Since the reference image already includes rich appearance information, the key is to design better motion representations and effectively guide the motion content generation process.
@@
Motion in videos can be decomposed into camera motion and object motion. Object motion refers to dynamic movements relative to the world coordinate system, while camera motion arises from viewpoint changes. Although researchers have made great progress in designing various motion representations to obtain control signals for camera or object, achieving collaborative camera and object motion control with fine control granularity remains non-trivial. For camera motion control, some methods directly use camera intrinsic and extrinsic parameters [37] or convert them into Plücker embeddings [41, 9, 51] as high-level control signals, while others reconstruct point cloud to render incomplete frames [3, 47, 11] or use dense optical flow [25, 17] as more concrete control signals. For object motion control, recent works usually follow an interactive setting that allows users to draw point trajectories on the reference frame. These methods convert sparse point trajectories into optical flow [37, 18, 49, 17, 25] or reorganize point embeddings [29, 38, 5, 4] as control signals. Nevertheless, works [17, 4, 37, 5] focusing on collaborative motion control still suffer from conflict control issues and insufficient control granularity.
@@
The key observations of this paper are three-fold: 1) Motion representations based on pixel offsets such as optical flow are flexible for in-the-plane motions, while the lack of 3D awareness leads to conflict control issues with simultaneous camera and object motions. 2) Reconstruction-based motion representations such as point clouds show better spatial consistency, making them effective for camera motion control but inflexible to support fine-grained object motion control. 3) For controllable visual content generation tasks that maintain basic spatial structures, spatially aligned control signals are easier for alignment than high-level ones. Based on these observations, an effective motion representation for fine-grained collaborative motion control should have the following properties. First, the motion representation is equipped with 3D awareness to disentangle local components from the reference image. Second, the motion representation is spatially consistent for stable camera motion control and flexible enough for fine-grained object motion control. Third, the motion control signals derived from the motion representation are spatially aligned with the generated frames to enhance their connections.
@@
Figure 2: 3D-aware motion representation for simplifying the 3D scene corresponding to the reference image, which consists of several unit spheres denoting key object parts and a world envelope accommodating them. We manipulate those spheres and observe the scene based on user intentions to transform camera and object motions into unified and consistent visual changes.
@@
To this end, we introduce 3D-aware motion representation. In controllable image animation applications, users generally have clear intentions about how objects move and how to observe the scene. The intuitive approach is to reconstruct the 3D scene based on the reference image, drive objects according to user intentions, and render the scene from the expected viewpoints. Considering the significant effort required to reconstruct everything in the scene, we notice that human visual attention is predominantly drawn to moving parts, while the remaining parts are generally perceived as a whole [6]. Therefore, as shown in Fig. 2, we simplify and summarize the 3D scene as a combination of several unit spheres representing key object parts (most moving parts and a few static parts) and a vast world envelope akin to The Truman Show accommodating them to demonstrate how the scene evolves as the observer’s viewpoint changes. We then render the instantiated 3D scene to obtain perception results that reflect camera and object motions and necessary spatial relations.
@@
Based on 3D-aware motion representation, we propose an image animation framework to achieve fine-grained collaborative motion control. The framework takes a reference image and the perception results of the corresponding 3D-aware motion representation as motion control signals, which are spatially aligned with the generated frames. Specifically, we apply two lightweight encoders to encode camera and object control signals separately to avoid RGB-level interference, and then merge them to form the final motion control signals. We employ a U-Net-based diffusion model and inject both the reference image and motion control signals, enabling motion-controllable image animation. A data curation pipeline is introduced to construct 3D-aware motion representations from in-the-wild videos. Moreover, we propose a three-stage training strategy to balance camera and object motion control during training.
@@
  • We introduce 3D-aware motion representation to facilitate flexible and fine-grained collaborative motion control, which transforms camera and object motions into unified and consistent visual changes and is orthogonal to specific base model architectures.
  • We propose an image animation framework to incorporate 3D-aware motion representation as spatially aligned motion control signals, which achieve collaborative motion control precisely and harmoniously.
  • The proposed approach can support various motion-related applications via translating user intentions, such as motion generation, motion clone, motion transfer, and motion editing.
@@ 2 Related Work
Motion-controllable image animation has made significant progress recently. Since reference images have greatly limited the degree of freedom in generating contents, achieving fine-grained collaborative motion control is even harder for image animation than text-based video generation.
@@
Camera Motion Control. Works for camera motion control [37, 41, 51, 9, 3, 47, 8, 31] aim to generate novel view images with spatial consistency. In computer graphics, camera information is typically represented using a set of intrinsic and extrinsic parameters. Methods like MotionCtrl [37] directly repeat these flattened camera parameters to image size as control signals. CameraCtrl [9] and CamCo [41] transform the camera parameters into Plücker coordinates as pixel-wise embeddings, which is more discriminative. ViewCrafter [47] and I2V-Camera [3] lift the reference image to 3D and render the point cloud to generate incomplete frames. SparseCtrl [8] and DimensionX [31] generate a specific set of training data for each type of camera motion and train individual LoRA models [12] tailored for each camera motion pattern. Concurrent work TrajectoryAttention [40] uses dense pixel trajectories to enhance temporal consistency, which is effective for camera motion control but hard to further support object motion control. Compared to high-level camera motion representations, our 3D-aware motion representation embodies camera motion information via visual changes of the world envelope, which is more intuitive and effective.
@@
Object Motion Control. Previous works usually follow the drag-based interaction setting for object motion control [46, 18, 29, 25, 49, 38], which conveys user intentions via sparse point trajectories. Methods such as DragNUWA [46], ImageConductor [18], Motion-I2V [29] and MOFA-Video [25] employ optical flow as an intermediate motion representation. The former two methods transform point trajectories into sparse optical flow as control signals, while the latter two methods transform sparse point trajectories into dense optical flows via off-the-shelf or pretrained sparse-to-dense flow generation models. Tora [49] enhances point trajectories into a trajectory map and employs 3D VAE to encode it. DragAnyhting [38] obtains features for each pixel via DDIM inversion [30] and reorganizes those features according to segmentation masks and point trajectories while being less effective in handling motions containing shape variations. Although the above methods support camera control using point trajectories, they suffer from conflict control issues in local regions with simultaneous camera and object motions. Instead of modifying 2D motion representation, we employ a simplified 3D scene and use its inherent perception results as control signals to avoid control conflicts. Due to this design, diffusion models can handle both local object motion and global camera motion simultaneously.
@@
Collaborative Motion Control. Only a few concurrent works [17, 5, 4] support collaborative motion control. AnimateAnything [17] applies two-stage generation, one for transforming camera and object control into optical flow representation, the other for generating videos based on that. However, they represent camera motion as Plücker embeddings, which is heterogeneous to optical flow and may lead to imprecise control. MotionPrompting [5] employs dense point embeddings constructed from 2D point tracking results as motion control signals, while such representation is not friendly enough for complicated motion control. I2VControl [4] also employs dense point embeddings constructed via point transform and projection, while the granularity of object motion control is limited to object segmentation masks. In contrast, we propose Perception-as-Control, which takes motion control signals derived from efficiently constructed 3D-aware motion representation and achieves fine-grained collaborative motion control.
@@ 3 Perception-as-Control
Figure 3: Overview. In our data curation pipeline, an in-the-wild video is taken as input, and off-the-shelf algorithms are applied to obtain camera poses and 3D point locations. Based on the results, 3D-aware motion representation can be constructed and rendered to derive camera and object control signals. During training, those control signals are fed into separate encoders to avoid RGB-level interference and merged afterward. A three-stage training strategy is introduced to achieve fine-grained collaborative motion control. During inference, user intentions are first interpreted as 3D-aware motion representation, enabling our framework to support various motion-related applications.
As shown in Fig. 3, our approach consists of three primary parts: a data curation pipeline for constructing 3D-aware motion representation from in-the-wild videos, an image animation framework for motion-controllable video generation, and a three-stage training strategy to facilitate fine-grained collaborative motion control.
@@ 3.1 Preliminaries
Video Diffusion Model (VDM) [2, 7, 44] adopts the foundational principles of image diffusion probabilistic models and models the temporal dependencies between frames, resulting in videos with consistent visual content and coherent motion. Let $x_{0}\in\mathbb{R}^{L\times H\times W\times C}$ denote the initial video latent, where $L$ is the number of frames. The forward process is defined as a chain that progressively adds Gaussian noise to the original video, as shown in Eq. 1:
$x_{t}=\sqrt{\hat{\alpha_{t}}x_{t-1}}+\sqrt{{1-\hat{\alpha_{t}}}\epsilon},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $t\in\{1,\dots,T\}$ denotes diffusion timestep, $\hat{\alpha_{t}}$ regulates the intensity of noise added at each $t$, and $\epsilon$ is drawn from standard Gaussian noise. The optimization objective is to predict the probability distribution $p(x_{t-1}|x_{t})$ using a parameterized network $\theta$. The loss function is in Eq. 2:
@@
where $\mathcal{C}$ is an optional parameter denoting the conditions, such as reference images and motion control signals.
@@
For training purposes, pairs of video clip and the corresponding 3D-aware motion representation are required. Unfortunately, commonly used large-scale video datasets predominantly consist of in-the-wild videos lacking camera settings and scene information. Reconstructing a delicate 3D scene is computationally expensive and the scale ambiguity issue leads to non-negligible reconstruction errors. Alternatively, constructing the simplified 3D scene directly via appropriate approximations can bypass these problems and is readily achievable even for in-the-wild videos. Thus, we introduce our data curation pipeline, using off-the-shelf algorithms to construct 3D-aware motion representation for any collectible video.
@@
SceneCut Detection and Motion Filtering. First, we apply common preprocessing steps to extract video clips containing continuous visual contents. We employ PySceneDetect [33] to split videos into video clips. Then, we use RAFT [32] to estimate optical flow for each video clip and use the Frobenius norm as a motion score to filter out video clips with non-obvious motion (motion scores lower than $30^{\rm th}$ percentile), the same as previous methods [18].
@@
3D Point Tracking for Local Object Motion. Key object parts refer to not only moving object parts but also stationary ones, which are critical for revealing spatial relations. They are simplified as unit spheres in the 3D scene, and the number of these spheres is defined according to the control granularity flexibly. We require 3D locations of key object parts at each moment to set these spheres initially and manipulate their movements. Therefore, we employ SpaTracker [39] to achieve 3D point tracking in each video clip. Note that combining algorithms of 2D point tracking [15] and depth estimation [26, 42, 43] also helps. In our implementation, we track $25\times 25$ grid points for each video clip, resulting in 3D point tracking results $\mathbf{P}=[\mathbf{P}^{\rm x},\mathbf{P}^{\rm y},\mathbf{P}^{\rm z}]^{\mathsf{T}}\in\mathbb{R}^{L\times N\times 3}$ ($L$ is the number of frames and $N$ is the number of tracked points).
@@
Camera Pose Estimation for Global Camera Motion. We employ a world envelope to mark the remaining parts, which is stationary relative to the world coordinate system. To transform global camera motion into visual changes, we can use off-the-shelf algorithms [36, 50] to recover camera trajectories from a sequence of continuous frames and render the world envelope. Despite the scale ambiguity problem, our world envelope, which serves as a scene marker, only requires coarse camera poses instead of extremely precise ones. In our implementation, we employ TartanVO [36] for this purpose, since it is fast and generalizes well across various cameras. Camera intrinsic parameters are set to $\mathbf{K}=[W,0,W//2;0,H,H//2;0,0,1]$, where $W$ and $H$ are the width and height of the first frame. For each video clip, we obtain the estimated camera poses $\mathbf{E}=\{\mathbf{E}_{l}|l\in\{1,2\dots,L\}\}$, where $L$ is the number of frames, $\mathbf{E}_{l}=[\mathbf{R}\in\mathbb{R}^{3\times 3},\mathbf{t}\in\mathbb{R}^{1\times 3}]$ are camera extrinsic parameters at frame $l$.
@@
Perceptual 3D Scene Rendering. Once we have obtained the 3D point tracking results $P$ and the estimated camera poses $E$, we construct our 3D-aware motion representation and render it as motion control signals.
@@
First, we set $N$ unit spheres with different colors representing key object parts. These spheres are rendered on the pixel plane according to a specific camera pose and appear as circles that are larger near and smaller far away in the rendering images under perspective projection. Since the estimated depth is not precise enough in some cases, we normalize point depths $\mathbf{P}^{\rm z}$ as $\mathbf{P}^{\rm d}$ for each video clip to only focus on their relative distance. After that, we set the projected circle radius of the $n^{\rm th}$ sphere to $1-\mathbf{P}^{\rm d}_{l,n}$ ($\mathbf{P}^{\rm d}_{l,n}$ is the normalized depth of the $n^{\rm th}$ sphere at frame $l$). For convenience, we generate a continuous color map where each pixel has a distinct color. We then set the color of each sphere to the corresponding color of its projection center in the first frame. Finally, these spheres are projected and rendered as colorful circles with varying radii.
@@
Then, we construct a cube centered at the origin with the side length of $z_{\rm far}$ ($z_{\rm far}$ is the far clipping plane of a viewing frustum) as an implementation of our world envelope. We apply checkerboard textures to increase the recognizability of the world envelope, which helps demonstrate how the scene evolves as the viewpoint changes. We align the camera coordinate system to the world coordinate system for the first frame and render the subsequent frames according to the relative camera pose sequence $E$, without loss of generality. Note that the implementation of the world envelope is not limited to a cube, and its textures can vary beyond the checkerboard style. To avoid RGB-level interference, we render unit spheres and world envelope separately as two spatial layers, acting as object and camera control signals respectively.